### v2

# Splitting criteria for ordinal decision trees: an experimental study

Abstract
Ordinal Classification (OC) is a machine learning field that addresses classification tasks where the labels exhibit a natural order. Unlike nominal classification, which treats all classes as equally distinct, OC takes the ordinal relationship into account, producing more accurate and relevant results. This is particularly critical in applications where the magnitude of classification errors has implications. Despite this, OC problems are often tackled using nominal methods, leading to suboptimal solutions. Although decision trees are one of the most popular classification approaches, ordinal tree-based approaches have received less attention when compared to other classifiers. This work conducts an experimental study of tree-based methodologies specifically designed to capture ordinal relationships. A comprehensive survey of ordinal splitting criteria is provided, standardising the notations used in the literature for clarity. Three ordinal splitting criteria, Ordinal Gini (OGini), Weighted Information Gain (WIG), and Ranking Impurity (RI), are compared to the nominal counterparts of the first two (Gini and information gain), by incorporating them into a decision tree classifier. An extensive repository considering $45$ publicly available OC datasets is presented, supporting the first experimental comparison of ordinal and nominal splitting criteria using well-known OC evaluation metrics. Statistical analysis of the results highlights OGini as the most effective ordinal splitting criterion to date. Source code, datasets, and results are made available to the research community.
keywords:
Ordinal classification , ordinal regression , ordinal trees , impurity measures , splitting criteria , ordinal Gini , ordinal information gain , ranking impurity
journal: Pattern Recognition
[…]
[uco]organization=Department of Computer Science and Numerical Analysis, Universidad de Córdoba, addressline=Campus de Rabanales, Ctra. N-IVa, Km. 396, city=Córdoba, postcode=14071, country=Spain
[…]
[uloyola]organization=Department of Quantitative Methods, Universidad Loyola Andalucía, addressline=C. Escritor Castilla Aguayo, 4, city=Córdoba, postcode=14004, country=Spain
1 Introduction
Machine Learning (ML) is one of the most important areas of artificial intelligence, with its primary goal being to enable models to learn patterns and make predictions on previously unseen data. Among the various ML paradigms, nominal classification plays an important role, in which the focus is on assigning discrete labels to instances according to learned representations from data. This paradigm can be found in a wide range of applications, ranging from specific medical procedures [1], such as donor-recipient matching in liver transplantation according to the survival time endpoints, to facing illegal trade [2], such as detecting levels of forged alcohol in spirit bottles. Despite their differences, these examples share a common characteristic: the set of labels exhibits a natural order relationship. In these cases, there is a field of ML, known as Ordinal Classification (OC) better suited for addressing this type of problems in which the labels are arranged following a given scale, and the distance between the labels is unknown. This field is also often referred to as ordinal regression.
As shown in the previous examples, OC problems are sometimes addressed by a standard nominal approach, obtaining non-optimal solutions. The main reason is that nominal methods disregard the order information of classes. In many real world applications, this information is essential, since the magnitude of the classification error could result in critical consequences. For example, in computed tomographic images of lung nodules [3], each image is assigned a label ranging from $1$ to $5$ representing the malignancy score. A value of $1$ indicates an low likelihood of malignancy, $3$ represents indeterminate malignancy, and $5$ reflects a high probability of malignancy. It is clear that classifying an image as indeterminate (scores $2$ or $3$) or even as non-malignant (score $1$), when it is actually highly malignant (score $5$), should be more penalised than classifying a non-malignant image (score $1$) as indeterminate (scores $2$ or $3$). This example perfectly reflects that the primary objective of OC is to bear in mind this inherent ordering to develop more accurate models.
Figure 1: Synthetic OC problem. Samples are coloured based on the class to which they belong. More information about this problem can be found in [4].
A synthetic example of an OC problem is illustrated in Figure 1, where samples are represented in a 2D feature space. Each sample is coloured according to its assigned class, with five different classes labelled from $C_{1}$ (red) to $C_{5}$ (green). The ordinal nature of the problem is evident, as the distance in the feature space increases with the ordinal difference between the classes. The boundaries separating the classes are shown, highlighting the gradual transition between adjacent classes and the greater distinction between classes further apart on the ordinal scale.
OC problems can be found in a wide variety of fields, including aforementioned medical research [5, 6, 7, 8, 9], automation science and engineering [10], wave height prediction [11], credit rating [12], face recognition [13, 14, 15], age estimation [16], image classification [17], wind speed prediction [18, 19], manufacturing [20], text classification [21], and transportation [22], among others. These studies exemplify the use of specifically designed OC models, which leverage the ordinal information to enhance performance compared to their nominal counterparts.
The OC paradigm has considerably evolved in recent years, thanks to substantial advances in supervised learning methodologies [23]. This includes contributions ranging from neural network models, like the one proposed in [24], which also addresses the imbalance problem typical of ordinal tasks by minimizing a Bayesian cost, to deep ordinal models [25], soft-labelling methods [26, 27] and ordinal representation learning techniques [28]. Tree-based models for OC have also been a hot topic in the last decade [29, 30, 31, 32]. One of the primary advantages of tree-based approaches is their ability to produce interpretable models that facilitate the understanding of complex and non monotonic dependencies between explanatory attributes and the dependent variable, also referred as output variable. For instance, the simplest tree-based models, the Decision Trees (DTs), can generate easy-to-interpret rules without compromising performance. Another strength of DTs is their versatility, as they can handle continuous, discrete and categorical inputs while being invariant under strictly monotone transformations of individual inputs. In addition, an internal feature selection is performed as an integral part of the training process.
The idea behind a DT involves recursively partitioning the predictor space, also known as input space, into a set of rectangles according to a splitting criterion. Subsequently, a simple model, such as a constant, is fitted to each rectangle. In standard nominal classification, the most widely used approaches are CART [33], along with other variants such as ID3 [34] and C4.5 [35].
This splitting criterion plays a key role in the learning process, as it determines how the predictor space is partitioned at each step, thereby guiding the construction of the tree. The splitting criteria used to assess the quality of a split is based on the concept of impurity, which quantifies the homogeneity of the labels at a given node. The objective during splitting is to minimise the impurity. Consequently, there is a strong relationship between designing effective splitting criteria and developing appropriate impurity measures. In nominal classification, the Gini-index [33] and Shannon-entropy [36] are the most used impurity measures, forming the basis of the well-known Gini and Information Gain (IG) splitting criteria.
Specifically, the Gini-index impurity measure computes the expected loss on all classes due to misclassification errors, whereas the Shannon-entropy represents the uncertainty of a random variable outcome. However, both Gini-index and Shannon-entropy, to be minimised, do not take into account the ordinal properties of the dependent variable. For instance, lets suppose the following two nodes representing the distribution of $20$ observations into $4$ ordered classes: A) $(10,0,0,10)$, and B) $(10,10,0,0)$. Considering these two distributions, we can observe that both Gini-index and Shannon-entropy result in the same value ($0.5$ and $1.0$, respectively). However, from an ordinal perspective, option B is preferable over option A, as it better separates the classes following an ordinal scale. This example is graphically illustrated in Figure 2, where the values for ordinal impurity measures are also presented (detailed along with their associated splitting criterion in Section 3). This observation extends to the splitting criteria, to be maximised, defined by these impurity measures. Figure 3 illustrates two different splits of a given node, for which the nominal splitting criteria achieve the same value regardless the ordinality considered in the sub-nodes.
The primary goal of this work is to address this limitation by analysing tree-based methods specifically designed to incorporate and leverage the ordinal information of the datasets. These methods aim to improve performance while maintaining the inherent structure of ordinal data, ensuring more relevant results. Other contributions of this work are the following:
  1. 1.
    To provide a comprehensive description and survey of existing splitting criteria for tree-based approaches in OC.
  2. 2.
    To standardise and unify the diverse notations previously used in the literature for existing ordinal splitting criteria, enabling consistency and clarity.
  3. 3.
    To present an extensive OC repository including a total of $45$ publicly available datasets to facilitate research and benchmarking.
  4. 4.
    To perform the first experimental study comparing the performance of several splitting criteria with their nominal counterparts. This comparison is conducted using three well-known performance metrics: Mean Absolute Error (MAE), Quadratic Weighted Kappa (QWK), and Ranked Probability Score (RPS).
  5. 5.
    To provide practitioners with a comprehensive Python implementation of various splitting criteria and the DT algorithm [^1] https://github.com/ayrna/decision-trees-from-scratch. We have developed a new framework that serves as a versatile platform for testing and validating novel approaches and splitting criteria in a plug-and-play way. Furthermore, the OC dataset repository and the experimental results are made available on the associated webpage [^2] https://www.uco.es/ayrna/grupos/ordinal-trees.
The rest of the paper is organised as follows: Section 2 reviews commonly used impurity measures (Section 2.1) and main OC tree-based approaches (Section 2.2); Section 3 presents the nominal and ordinal impurity measures and their corresponding splitting criteria; Section 4 describes the experimental study conducted; the results achieved are included and statistically analysed in Section 5; and finally, Section 6 closes this work with some final remarks and future research directions.
2 Literature review
This section examines the most common impurity measures used in DT for nominal classification, discussing their limitations in capturing ordinal relationships and highlighting the need for more appropriate impurity measures for OC problems. Then, a review of the main tree-based methodologies for OC is provided, addressing a gap in the literature as no prior surveys have specifically explored this area.
2.1 Impurity measures
An impurity measure is a metric that estimates the probability of misclassifying a randomly chosen sample from a labelled set. This can also be observed as the quantification of the heterogeneity of a set of labelled patterns with respect to their labels. A higher impurity value indicates greater heterogeneity, while a lower value reflects greater homogeneity. The Gini-index and Shannon-entropy are the two most used measures in this regard.
A generalisation of the Gini-index impurity specifically designed for DT-based methodologies was proposed in [33], expressed as the misclassification cost of assigning a sample to a wrong class. An extension of the Gini-index appropriate for OC was presented in [37], known as Ordinal Gini-index (OGini-index). The idea behind OGini-index is to employ the cumulative frequency of each class instead of the relative frequency to measure the heterogeneity.
On the other hand, the Shannon-entropy ($H$) [36] is calculated based on the probabilities of each class. Several works have extended the Shannon-entropy to take into account ordinal information of the classes. However, most of them assume monotonicity not only of the output but also of the input variables [38, 39, 40]. Recently, a variant of Shannon-entropy was introduced in [41], referred to as Weighted Entropy ($H_{w}$). Unlike traditional entropy measures, this impurity measure does not depend on the assumption of monotonic behaviour and was specifically adapted for ordinal DTs in [42]. The weighting mechanism in this work was designed to account for the ordinal nature of the dataset by calculating the distances between classes relative to the mode class, which represents the most probable class. Furthermore, $H_{w}$ was successfully applied to Random Forest (RF) and AdaBoost algorithms [7, 43], extending its utility to ensemble methods. Building on these advancements, [32] provided a generalisation of $H_{w}$, making it applicable to any DT-based methodology.
Finally, there exists an impurity measure known as the Ranking Impurity (RI) [44], which, unlike previous impurity measures, was not derived from a nominal counterpart. The RI can be interpreted as the maximum potential number of miss-ranked pairs in the set. Here, the miss-ranked pairs are weighted by the difference between the assigned scores.
As observed, most of the aforementioned ordinal impurity measures depend on the assignment of numerical scores to the classes, thereby using a higher scale level. Specifically, the $H_{w}$ and RI explicitly incorporate these assigned scores into their definitions. In contrast, OGini-index is score-free as its definition is based on the cumulative frequency of each class. While the assignment of scores may be justified in certain scenarios, particularly when ordinal labels are built from continuous variables by discretisation, it is rather artificial and arbitrary in the context of genuine ordinal data. For instance, when the response variable represents ordered levels of disease severity, assigning numerical scores might not align with the natural characteristics of the data. In this work, we use the three previous ordinal impurity measures to build different splitting criteria, which are thoroughly defined and mathematically formalised in Section 3.
2.2 Ordinal classification trees
The first review on ordinal classification, to the best of the authors’ knowledge, was presented in [45], where a general taxonomy was proposed based on how models exploit ordinal information. In this survey, the approaches were categorised into three distinct groups: 1) naive approaches, which involve standard ML algorithms designed for nominal classification or regression; 2) ordinal binary decomposition techniques, where an ordinal problem with $Q$ classes is decomposed into $Q$ binary classification problems; and 3) threshold-based techniques, which assume an underlying continuous variable that is modelled using a set of thresholds that map intervals on the real line to the corresponding ordinal labels.
A more recent review was presented in [31] with an up-to-date taxonomy covering novel approaches and most prominent methodologies. In this case, the taxonomy includes the following 4 groups: 1) basic OC models, which include cumulative models, sequential models or adjacent categories models, among others; 2) models with more complex parametrisations, which extend previous basic models by including category-specific effects or capture intricate relationships between inputs; 3) hierarchically structured models, that are flexible approaches for ordinal data that exploit the inherent structure of response categories by sequentially modelling grouped and subdivided categories; and 4) finite mixture models, which combine content-driven labels with uncertainty or response patterns, using probabilistic mixtures to capture heterogeneity.
However, some approaches do not properly fit into the categories defined by either taxonomy, such as the tree-based approaches, in which a number of methods have been proposed in recent years. Beyond the approaches detailed in Section 2.1, additional methods have been introduced in the literature, such as the one proposed in [4], which is based on the concept of consistency, a global property that characterises the relationship between different decision regions within the input space. This property was integrated into the design of the learning algorithm, where it guides the node expansion process to ensure alignment with the ordinal structure of the data.
Furthermore, building upon the RF methodology, three notable approaches have been introduced in the literature. The first one, known as Ordinal Forest [46], leverages the latent variable underlying the observed ordinal output. This method combines the strengths of ensemble models with the threshold-based approach to reduce the bias and variance associated with individual ordinal classifiers. The second study, presented by Janitza et al. in [30], extends RF by incorporating conditional inference trees, offering a robust framework for OC tasks. Finally, the third work proposed a variant of ordinal trees (also used in RF) based on binary models that are implicitly used in parametric OC [47].
3 Methodology
Let’s define a dataset $D$ with $N$ patterns as $D=\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{i},y_{i}),\ldots,(\mathbf{x}_{N},y_{N})\}$ where $\mathbf{x}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{K}$ represents a vector of $K$ independent variables (or features), and $y_{i}\in\mathcal{Y}=\{C_{1},C_{2},\ldots,C_{q},\ldots,C_{Q}\}$ is the label assigned to input pattern $\mathbf{x}_{i}$, i.e. the dependent variable. Since this work focuses on ordinal classification problems, it is assumed that $C_{1}\prec C_{2}\prec\ldots\prec C_{Q}$, where $\prec$ represents the order relationship. The standard approach in ordinal problems is to assign a numerical value to each category $C_{q}\in\mathcal{Y}$. This assignment is defined by a function $v$, normally set as $v(C_{q})=q$, satisfying that $v(C_{1})<v(C_{2})<\ldots<v(C_{q})<\ldots<v(C_{Q})$. An analogy to this ordering can be found in regression problems, where the $\prec$ is replaced by the standard $<$ operator, and $\mathcal{Y}=\mathbb{R}$. However, unlike regression, in ordinal classification, the exact distance between classes is unknown. Nevertheless, it is assumed that different penalisation should be given to different classification errors, such that misclassifying a pattern into an adjacent class incurs a smaller penalty compared to misclassifying it into a class farther away on the ordinal scale.
In general terms, a DT is constructed by recursively partitioning the input dataset, so that patterns belonging to the same class are grouped together and separated as much as possible from patterns belonging to different classes. At each node $m$ of a DT, we work with a subset $D_{m}\subseteq D$. For the root node ($m=0$), $D_{m}=D$. During the subsequent steps of the growth phase, at each node $m$, the goal is to partition $D_{m}$ into two subsets $D_{m}^{l}$ (left sub-node) and $D_{m}^{r}$ (right sub-node). This partition is made by a split $\theta(k,t_{k})$, where $k$ denotes one of the $K$ features, and $t_{k}$, is a real-valued threshold that specifies the splitting point for the feature $k$. Then, the two subsets $D_{m}^{l}$ and $D_{m}^{r}$ are defined as follows:
$D_{m}^{l}=\{(\mathbf{x}_{i},y_{i})\mid x_{i_{k}}<t_{k}\},$ (1)
$D_{m}^{r}=\{(\mathbf{x}_{i},y_{i})\mid x_{i_{k}}\geq t_{k}\},$
where $x_{i_{k}}$ is the value for the $k$-th feature of the $i$-th pattern. According to [33], the optimal split $\theta(k^{*},t_{k}^{*})$ is the one maximising the real-valued function $\phi(D_{m},\theta)\in\mathbb{R}$, which is commonly known as the splitting criterion. In general, the goal of the $\phi$ function is to quantify the impurity decrease that results from splitting a node, hence, it has to be maximised. As it was previously defined, the impurity is a measure of the heterogeneity with respect to the output labels $\mathcal{Y}_{m}=\{y_{i}\in\mathcal{Y}:y_{i}\in D_{m}\}$ within a node $m$. Since the development of DTs, numerous splitting criteria have been introduced in the literature. Among these, the Gini and Information Gain (IG), derived from the Gini-index and Shannon-Entropy impurity measures, respectively, have demonstrated the strongest performance [48].
3.1 Gini
The Gini-index is defined as a function $G$ that quantifies the probability of misclassifying a sample randomly drawn from the node, assuming it was labelled based on the class distribution of that node. The Gini-index of a given set $D_{m}$ is defined as:
$G(D_{m})=1-\sum_{q=1}^{Q}{\frac{N_{q}(D_{m})}{N(D_{m})}},$ (2)
where $N_{q}(D_{m})$ represents the number of patterns belonging to category $C_{q}$ in $D_{m}$, and $N(D_{m})$ is the total number of patterns in $D_{m}$. For simplicity, from now on, $\hat{p}_{q|m}=\frac{N_{q}(D_{m})}{N(D_{m})}$, i.e. the relative frequency of class $C_{q}$ at the $m$ node.
Therefore, the splitting criterion based on the Gini-index ($\phi_{Gini}$, for simplicity, denoted as Gini) measures the impurity decrease caused by splitting a node $m$ into its left and right child nodes as follows:
$\phi_{\text{Gini}}(D_{m},\theta)=G(D_{m})-(\hat{p}_{l}\ G(D_{m}^{l})+\hat{p}_{r}\ G(D_{m}^{r})),$ (3)
where $\hat{p}_{l}=\frac{N(D_{m}^{l})}{N(D_{m})}$ and $\hat{p}_{r}=\frac{N(D_{m}^{r})}{N(D_{m})}$ are the proportion of samples in the left and right nodes, respectively.
The Gini splitting criterion is highly effective for standard classification tasks. However, it disregards the ordinal relationship among classes. As a result, misclassifying a sample always incurs the same amount of loss, ignoring the distance in the ordinal scale between the observed and the predicted classes.
3.2 Ordinal Gini (OGini)
For ordinal classification, a variant of the Gini-index impurity measure, known as Ordinal Gini-index (OGini-index), was proposed in [37], defined as a function $OG$. The idea was to measure the heterogeneity between the two sub-nodes using the dissimilarity between the cumulative frequencies:
$OG(D_{m})=1-\sum_{q=1}^{Q}{\hat{c}_{q}},$ (4)
where $\hat{c}_{q}=\sum_{j=1}^{q}p_{j}$ is the cumulative frequency of class $C_{q}$ in $D_{m}$. Then, the splitting criterion based on $OG$ ($\phi_{OGini}$, for simplicity, denoted as OGini) is expressed as follows:
$\phi_{\text{OGini}}(D_{m},\theta)=OG(D_{m})-(\hat{p}_{l}\ OG(D_{m}^{l})+\hat{p}_{r}\ OG(D_{m}^{r})).$ (5)
3.3 Information Gain (IG)
The IG splitting criterion is based on the Shannon-entropy impurity measure [36]. In a DT setup, entropy, $H$, quantifies the uncertainty or disorder in a given node, $m$. If the node is perfectly homogeneous (all instances belong to the same class), $H=0$ because there is no uncertainty. On the other hand, if the node is equally split between classes, $H$ is at its maximum. Hence, $H$ at a node $m$ is defined as follows:
$H(D_{m})=-\sum_{q=1}^{Q}\left[\hat{p}_{q|m}\ log(\hat{p}_{q|m})\right].$ (6)
Therefore, the IG splitting criterion, $\phi_{IG}$, quantifies the reduction in $H$ caused by splitting a node as follows:
$\phi_{\text{IG}}(D_{m},\theta)=H(D_{m})-(\hat{p}_{l}\ H(D_{m}^{l})+\hat{p}_{r}\ H(D_{m}^{r})).$ (7)
3.4 Weighted Information Gain (WIG)
The standard entropy measure, $H$, does not take into account the ordinal properties of the target variable $\mathcal{Y}$. Different variants of $H$ were proposed in [41], standing out the weighted entropy, $H_{w}$. This approach allocates weights to the different ordinal classes in such a way that the weight $w_{q}(D)$, assigned to the class $C_{q}$, is derived from the difference between the value of the class and the value of the mode class, i.e. the most probable class. This weight is defined as:
$w_{q}(D)=\frac{|v(C_{q})-v(C_{\text{mode}})|^{\propto}}{\sum^{Q}_{j=1}|v(C_{j})-v(C_{\text{mode}})|^{\propto}},$ (8)
where $C_{\text{mode}}$ is the mode class, i.e. the class with more instances in $D$, and $\propto$ is a normalisation parameter that balances the distribution of weights over the different classes. The introduction of weights is intended to consider the distribution of classes along the ordinal scale. Therefore, $H_{w}$, is defined as:
$H_{w}(D_{m})=-\sum_{q=1}^{Q}\left[w_{q}(D_{m})\ \hat{p}_{q|m}\ \log(\hat{p}_{q|m})\right].$ (9)
With the goal of considering the ordinal properties of the ordinal output, the Weighted Information Gain (WIG) splitting criterion, $\phi_{WIG}$, was introduced in [42], replacing $H$ of Equation 7 by its weighted version, $H_{w}$, giving birth to:
$\phi_{\text{WIG}}(D_{m},\theta)=H_{w}(D_{m})-(\hat{p}_{l}\ H_{w}(D_{m}^{l})+\hat{p}_{r}\ H_{w}(D_{m}^{r})).$ (10)
It is important to highlight that the computation of the weights $w_{q}(D)$ for each class depends on the scores assigned to the ordered categories of the response, i.e. the function $v$ under consideration.
Additionally, different variations of weighted entropy, $H_{w}$, can be found in [7, 43], replacing, for instance, the mode of the classes by the highest class value, or employing those weights in an entropy-based objective function used to build the tree.
3.5 Ranking Impurity (RI)
The Ranking Impurity (RI) was presented in [49], with the focus of measuring the potential maximum number of miss-classified samples based on the current node distribution. A higher impurity value indicates more potential for misalignment between the observed and predicted classes. The RI at node $m$ is defined as follows:
$RI(D_{m})=\sum_{q=1}^{Q}\sum_{j=1}^{q}\left[\beta(C_{q},C_{j})\ N_{q}(D_{m})\ N_{j}(D_{m})\right],$ (11)
where $\beta$ is a weighting function that penalises the impurity according to the ordinal scale. The introduction of $\beta$ generalises the expression presented in [49], the one considered in our experimentation, in which $\beta(C_{q},C_{j})=v(C_{q})-v(C_{j})$. The splitting criterion based on RI ($\phi_{RI}$, for simplicity, the splitting criterion also adopts the RI acronym) is defined as:
$\phi_{\text{RI}}(D_{m},\theta)=RI(D_{m})-(\hat{p}_{l}\ RI(D_{m}^{l})+\hat{p}_{r}\ RI(D_{m}^{r})).$ (12)
In summary, the Gini and OGini splitting criteria are based on the Gini-index and OGini-index impurity measures, respectively. The standard IG and its ordinal counterpart, WIG, are derived from the standard entropy, $H$, and its weighted variant, $H_{w}$, respectively. Then, the RI splitting criterion is directly associated with the RI impurity measure.
The way impurity measures and splitting criteria are computed is graphically illustrated in Figures 2 and 3. Regarding impurity measures, Figure 2 depicts two distinct ordinal distributions for a problem with four classes. Note that impurity measures are to be minimised. From the perspective of an ordinal scale, the upper distribution exhibits poorer separation, as distant classes are grouped together. Conversely, the lower distribution demonstrates better separation by grouping adjacent classes. Nominal impurity measures (Gini and standard entropy, $H$) fail to account for this distinction. However, the ordinal impurity measures (OGini, weighted entropy – $H_{w}$, and $RI$) explicitly consider these differences and consequently assign more favourable (lower) values to the distribution located below.
Figure 2: Representation of two different ordinal distributions for a problem with four classes, and associated values for the different impurity measures considered in this work.
Figure 3 illustrates two different node splits, highlighting the differences between ordinal and nominal splitting criteria. The values in the figure represent splitting criteria that are maximised during the growing procedure. On the left side of the figure, the split divides the patterns such that the left sub-node contains all patterns for $C_{1}$ and $C_{4}$, while the right sub-node contains patterns for $C_{2}$ and $C_{3}$. In this case, the left sub-node does not maintain an ordinal order, as the classes within it are not contiguous. Conversely, the split shown on the right adheres to an ordinal structure: the left sub-node includes patterns for $C_{1}$ and $C_{2}$, and the right sub-node contains patterns for $C_{3}$ and $C_{4}$, preserving the order relationship. This distinction is also reflected in the splitting criteria values presented below the nodes, where higher values are observed for the ordinal splitting criteria in the right split. Notably, nominal splitting criteria achieve the same values regardless of whether the splits maintain ordinality or not.
Figure 3: Comparison of two node splits based on nominal and ordinal splitting criteria. The left split does not preserve the ordinal relationship among classes. Conversely, the right split maintains the ordinal structure, achieving higher values for ordinal splitting criteria. Note that nominal splitting criteria remain unaffected by the ordinality of the splits.
4 Experimental setup
This section describes the experimental framework used to evaluate the differences in terms of performance for the various splitting criteria presented in Section 3. These splitting criteria include Gini, OGini, IG, WIG, and RI, which are utilised to build a DT classifier. Note that all other hyperparameters of the DT classifier remain fixed, with each DT classifier being named after the splitting criterion used during its training phase.
To conduct a comprehensive study, we evaluate each splitting criterion using a set of $45$ ordinal datasets collected from different OC works. Specifically, $41$ of them were obtained from [45], which is the the most extensive experimental study conducted to date in the OC field. The remaining $4$ datasets are well-established and frequently utilised [30]. A detailed overview of these datasets is provided in Table 1.
For consistency and fairness, each splitting criterion is evaluated on every dataset over $20$ independent runs per dataset, using different seeds. To ensure comparability, the same set of seeds is applied across all splitting criteria. In each run, the maximum depth of the trees is adjusted using a stratified $5$-fold crossvalidation procedure, optimised in terms of Mean Absolute Error (MAE), and selecting the best value from the set $\{3,5,8,16\}$. Moreover, the tree-growing method for the DT classifier is restricted to the one used in CART method [33]. In the special case of the WIG criterion, the normalisation parameter $\propto$ is set to $1$.
For the evaluation, we consider three different evaluation metrics that take the ordinality of the target variable into account: MAE, Quadratic Weighted Kappa (QWK), and Ranked Probability Score (RPS). The MAE is a widely utilised metric in OC. It measures the mean absolute difference between the observed and the predicted classes, considering the scores assigned by the function $v$. It is expressed as:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|v(y_{i})-\hat{y}_{i}|.$ (13)
The QWK is based on the kappa index [50], which measures the inter-rater agreement on classifying elements into a set of categories. The value of the QWK ranges from $-1$ to $1$, with $1$ being the best possible score. This metric incorporates a penalisation term that grows quadratically with the distance between the two predicted classes. In this way, the QWK is expressed as:
$\text{QWK}=1-\frac{\sum_{q,j}W_{q,j}O_{q,j}}{\sum_{q,j}W_{q,j}E_{q,j}},$ (14)
where $\mathbf{W}$, $\mathbf{O}$ and $\mathbf{E}$ are matrices of size $Q\times Q$. $\mathbf{O}$ is the confusion matrix. $\mathbf{E}$ represents the expectation to obtain the observed agreement by random chance, where each element $E_{q,j}$ is computed as $E_{q,j}=\frac{\sum_{i}\mathds{1}(y_{i}=C_{q})\sum_{i}\mathds{1}(\hat{y}_{i}=v(C_{j}))}{N}$. Here, $\hat{y}_{i}$ is the predicted category for the $i$-th instance. The value of the $W_{q,j}$ element of the quadratic weights matrix $W$ is computed as:
$W_{q,j}=\frac{(v(C_{q})-v(C_{j}))^{2}}{(Q-1)^{2}}.$ (15)
Finally, the RPS metric was introduced in [51] as a generalisation of the Brier score. It is a minimisation metric that measures the averaged difference between the predicted and the observed cumulative distribution of the ordinal classes:
$\text{RPS}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q=1}^{Q}(\hat{\pi}_{i}(q)-\mathds{1}(y_{i}\leq C_{q}))^{2},$ (16)
where $\hat{\pi}_{i}(q)$ denotes the estimated cumulative probability of observing class $C_{q}$ in the $i$-th sample, and $\mathds{1}(y_{i}\leq C_{q})=1$ when $y_{i}\leq C_{q}$ is true, and 00 otherwise.
Table 1: Characteristics of the datasets used in this work. #Train and #Test stand for the number of training and testing patterns, respectively, $Q$ indicates the number of classes, and $K$ is the number of input features.
Original OC problems Discretised regression problems
Dataset #Train #Test $Q$ $K$ Dataset #Train #Test $Q$ $K$
ERA $750$ $250$ $9$ $4$ abalone $1000$ $3177$ $5$ $10$
ESL $366$ $122$ $9$ $4$ abalone10 $1000$ $3177$ $10$ $10$
LEV $750$ $250$ $5$ $4$ bank1-10 $50$ $8142$ $10$ $8$
SWD $750$ $250$ $4$ $10$ bank1-5 $50$ $8142$ $5$ $8$
automobile $153$ $52$ $6$ $71$ bank2-10 $75$ $8117$ $10$ $32$
balance-scale $468$ $157$ $3$ $4$ bank2-5 $75$ $8117$ $5$ $32$
bondrate $42$ $15$ $4$ $37$ calhousing-10 $150$ $20490$ $10$ $8$
car $1296$ $432$ $4$ $21$ calhousing-5 $150$ $20490$ $5$ $8$
contact-lenses $18$ $6$ $3$ $6$ census1-10 $175$ $22609$ $10$ $8$
eucalyptus $552$ $184$ $5$ $91$ census1-5 $175$ $22609$ $5$ $8$
mammoexp $276$ $136$ $3$ $5$ census2-10 $200$ $22584$ $10$ $16$
newthyroid $161$ $54$ $3$ $5$ census2-5 $200$ $22584$ $5$ $16$
nhanes $3499$ $1724$ $5$ $30$ computer1-10 $100$ $8092$ $10$ $12$
pasture $27$ $9$ $3$ $25$ computer1-5 $100$ $8092$ $5$ $12$
squash-stored $39$ $13$ $3$ $51$ computer2-10 $125$ $8067$ $10$ $21$
squash-unstored $39$ $13$ $3$ $52$ computer2-5 $125$ $8067$ $5$ $21$
support $489$ $241$ $5$ $19$ housing $300$ $206$ $5$ $13$
tae $113$ $38$ $3$ $54$ housing10 $300$ $206$ $10$ $13$
toy $225$ $75$ $5$ $2$ machine $150$ $59$ $5$ $6$
vlbw $115$ $57$ $9$ $19$ machine10 $150$ $59$ $10$ $6$
winequality-red $1199$ $400$ $6$ $11$ pyrim $50$ $24$ $5$ $26$
pyrim10 $50$ $24$ $10$ $26$
stock $600$ $350$ $5$ $9$
stock10 $600$ $350$ $10$ $9$
5 Results
The results obtained for the $20$ executions of each DT classifier using different splitting criteria are presented in Table 2, expressed as $\text{Mean}_{\text{STD}}$, where STD stands for the standard deviation. It can be observed that the ordinal splitting criteria, i.e. OGini, WIG, and RI, obtain better results for MAE and QWK than the nominal ones (Gini and IG). In terms of RPS, OGini and RI again outperform IG and Gini. However, in this case WIG fails to outperform the nominal metrics. It is important to remark that OGini obtains the best mean results for the three metrics, while the second best results are obtained by the WIG in MAE and QWK, and by the RI in terms of RPS.
Table 2: Mean and standard deviation (${}_{\text{STD}}$) for each method in terms of MAE, QWK and RPS. All datasets and splits are considered.
Method MAE ($\downarrow$) QWK ($\uparrow$) RPS ($\downarrow$)
Gini $0.875_{0.098}$ $0.642_{0.071}$ $0.745_{0.125}$
IG $0.880_{0.103}$ $0.641_{0.073}$ $0.749_{0.139}$
OGini $\mathbf{0.849_{0.094}}$ $\mathbf{0.658_{0.068}}$ $\mathbf{0.718_{0.129}}$
WIG $\mathit{0.869_{0.092}}$ $\mathit{0.650_{0.067}}$ $0.770_{0.123}$
RI $\mathit{0.869_{0.096}}$ $0.646_{0.068}$ $\mathit{0.719_{0.146}}$
The best result is highlighted in bold, whereas the
second best is in italics.
To analyse the statistical significance of the differences among the considered splitting criteria for MAE, QWK, and RPS performance metrics, the following statistical setup is used: 1) the non-parametric Kolmogorov-Smirnov normality test (K-S test) is firstly applied to verify if the results obtained for each metric follow a normal distribution; 2) given that the null hypothesis of normality is accepted by K-S test in our experiments, an ANalysis Of the VAriance (ANOVA) test [52, 53] is performed. Specifically, an ANOVA II is applied to analyse the effects of two qualitative factors (the dataset and the splitting criterion) on the quantitative response, i.e. the values obtained in terms of MAE, QWK, and RPS metrics; 3) if the null hypothesis of the ANOVA II test, indicating no significant differences between means, is rejected, a Tukey post-hoc test is finally conducted to identify the specific groups between which significant differences exist. A significance level of $\alpha=0.050$ is used for these statistical tests.
For further analysis on the scalability of these approaches, and following the previous statistical setup, two groups, based on the number of classes, $Q$, are considered. The motivation behind this lies in the assumption that the ordinal characteristics of a problem become more significant as $Q$ increases. As a result, the differences in performance between nominal and ordinal splitting methods are expected to be more pronounced in datasets with higher values of $Q$. Specifically, we first examine the datasets with $6$ or more classes ($Q\geq 6$) and then those with fewer than $6$ classes ($Q<6$). This threshold of $6$ is chosen as it approximately balances the groups: the first group consists of $17$ datasets, while the second group includes $28$ datasets.
5.1 Statistical analysis for datasets with $Q\geq 6$
In terms of MAE, the null hypothesis of normality of the K-S test is accepted for each of the $85$ samples of size $20$ ($17$ databases × $\times\ 5$ splitting criteria). Subsequently, the ANOVA II test is performed to test whether the methodology and dataset (factors) have any significant impact on the mean values obtained in each metric. The results indicate the existence of significant differences between the mean results obtained by the different methodologies (p-value $<0.001$). In view of these results, the Tukey post-hoc test is conducted. The results for this test are presented in Table 3, where the methods (splitting criteria) are grouped based on the statistical differences in their performance. Methods within the same group exhibit no significant differences, while those in different groups show statistically significant differences. As can be observed in the left part of Table 3 ($Q\geq 6$), group $1$ contains the best methodologies, OGini and RI, with no significant differences between them (p-value $=0.063$), and the group $3$ contains the worst methodologies, Gini and IG. This table also indicates significant differences between the pairs OGini-Gini, OGini-IG, WIG-IG, and RI-IG, always in favour of the ordinal methods. However, although RI presents a lower MAE, the difference against Gini is not significant (group $2$). Additionally, OGini stands as the best performing splitting criterion, obtaining significant differences against all the other criteria except for the RI.
Table 3: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of MAE.
$Q\geq 6$ $Q<6$
Group Rank Group Rank
Method $1$ $2$ $3$ Method $1$ $2$
OGini $1.393$ OGini $0.519$
RI $1.426$ $1.426$ Gini $0.526$ $0.526$
WIG $1.429$ IG $0.529$
Gini $1.454$ $1.454$ RI $0.531$
IG $1.464$ WIG $0.534$
p-value $0.063$ $0.153$ $0.945$ p-value $0.161$ $0.577$
For the QWK metric, the K-S test accepts the null hypothesis of normality, so the same procedure utilised for the MAE is conducted. The ANOVA II shows an overall statistically significant difference between the mean results across the splitting criteria (p-value $<0.001$). Therefore, the post-hoc Tukey test is conducted (Table 4). OGini and RI outperform both IG and Gini, and WIG surpasses IG. However, WIG shows no significant differences compared to Gini (p-value $=0.089$). Similarly as for the MAE metric, OGini stands as the best approach, having no significant differences against RI and WIG (p-value $=0.996$).
Table 4: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of QWK.
$Q\geq 6$ $Q<6$
Group Rank Group Rank
Method $1$ $2$ $3$ Method $1$
OGini $0.665$ OGini $0.654$
RI $0.656$ WIG $0.646$
WIG $0.653$ $0.653$ RI $0.643$
Gini $0.641$ $0.641$ IG $0.643$
IG $0.639$ Gini $0.642$
p-value $0.996$ $0.089$ $0.143$ p-value --
Finally, in terms of RPS, once the hypothesis of normality is accepted, the ANOVA II test demonstrate a significant difference between the mean results of the five splitting criteria (p-value $<0.001$). The Tukey test (see Table 5) indicates that there are no significant differences in mean RPS between RI, OGini, Gini and IG (p-value $=0.082$). On the contrary, WIG is observed to be significantly worse than RI, Ogini and Gini. Additionally, we can observe that RI obtains the best results.
Table 5: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of RPS.
$Q\geq 6$ $Q<6$
Group Rank Group Rank
Method $1$ $2$ Method $1$ $2$
RI $1.189$ OGini $0.437$
OGini $1.190$ Gini $0.440$
Gini $1.229$ RI $0.442$
IG $1.236$ $1.236$ IG $0.446$ $0.446$
WIG $1.284$ WIG $0.463$
p-value $0.082$ $0.066$ p-value $0.590$ $0.067$
As a general comment, by comparing left and right parts of Tables 3, 4 and 5, it is clear that differences are more evident in performance when the number of classes increases. This comes from the fact that the possible costs of misclassification increases, and ordinal splitting criteria can better exploit the ordinality of the target label and its relationship with the input data.
5.2 Statistical analysis for datasets with $Q<6$
Focusing on the MAE metric, the hypothesis of normality is accepted, then, the ANOVA II is applied, showing that there are significant differences between the mean values of the five splitting criteria (p-value$=0.049$). Thus, the post-hoc Tukey test is conducted. The results are shown in the right part of Table 3. It can be observed that OGini achieves the best mean results. However, there are no significant differences between OGini, RI, and WIG with respect to Gini and IG.
In terms of QWK, since the hypothesis of normality is accepted, the ANOVA II test is applied indicating that no significant differences are found with respect to the mean results obtained by the splitting criteria, achieving a p-value$=0.446$. In this case, the post-hoc Tukey test is not performed (in Table 4 all the splitting criteria are in the same group).
Finally, for the RPS metric the hypothesis of normality is accepted again. Then, the ANOVA II test is conducted, showing significant differences in the mean values between the different splitting criteria. Hence, the post-hoc Tukey test is performed. The results (see Table 5) indicate that OGini is the best splitting criteria in mean results, but no significant differences are found with respect to Gini and IG.
5.3 Analysis of the confusion matrices
In this subsection, the benefit of using ordinal splitting criteria compared to their nominal counterparts is illustrated using the census2-5 dataset. For this purpose, two DT classifiers are trained using the Gini and OGini splitting criteria, respectively, with a specific random seed. The resulting confusion matrices for each classifier are presented below, highlighting the differences in how the models classify the data. These matrices demonstrate the impact of considering ordinal information in the splitting process, as the OGini splitting criterion better accounts for the ordinal relationships between classes, leading to closer targets in the ordinal scale when compared to the Gini splitting criterion.
Table 6: Confusion matrices obtained in the test set by Gini and OGini in the census-2-5 dataset together with the QWK, MAE and RPS values.
Gini OGini
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$ $C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$
$C_{1}$ $2197$ $1289$ $438$ $342$ $251$ $2115$ $1475$ $842$ $77$ $8$
$C_{2}$ $1025$ $1440$ $859$ $678$ $515$ $932$ $1209$ $1895$ $414$ $67$
$C_{3}$ $475$ $1035$ $1115$ $1233$ $659$ $364$ $833$ $1905$ $1217$ $198$
$C_{4}$ $239$ $504$ $872$ $1922$ $980$ $151$ $407$ $1099$ $2292$ $568$
$C_{5}$ $193$ $331$ $360$ $1827$ $1805$ $79$ $236$ $529$ $2440$ $1232$
QWK = $0.534$ QWK = $0.648$
MAE = $0.948$ MAE = $0.793$
RPS = $0.948$ RPS = $0.571$
It can be observed that the use of the OGini splitting criterion significantly reduces errors in the extreme classes compared to the Gini criterion. For example, the number of instances where patterns from class $C_{1}$ are misclassified as class $C_{5}$ decreases dramatically from $251$ to $8$. Similarly, the number of instances where patterns from class $C_{5}$ are misclassified as class $C_{1}$ is reduced from $193$ to $79$. Moreover, considering the case of class $C_{3}$, which lies in the middle of the ordinal scale, the number of patterns misclassified as class $C_{1}$ and class $C_{5}$ is reduced from $475$ to $364$, and from $659$ to $198$, respectively. These improvements highlight the ability of ordinal splitting criteria to better respect the ordinal relationships between classes, particularly in cases where the misclassification involves distant classes on the ordinal scale.
6 Conclusion
This work provides an exhaustive survey of the ordinal splitting criteria proposed in the literature: Ordinal Gini (OGini), Weighted Information Gain (WIG), and Ranking Impurity (RI), where the first two are ordinal adaptations of the well-known Gini and Information Gain (IG). A unified notation has been established to clearly define the different splitting criteria, as well as the impurity measures associated to them: Gini-index and Shannon-entropy for the nominal splitting criteria, and OGini-index, weighted entropy, and RI, for the ordinal ones.
The methodologies have been empirically evaluated using a wide range of $45$ ordinal problems in terms of three different ordinal metrics: Mean Absolute Error (MAE), Quadratic Weighted Kappa (QWK), and Ranked Probability Score (RPS). The results presented in this work confirm the superiority of the ordinal splitting criteria over their nominal counterparts, in the mean values obtained for MAE, QWK and RPS, with OGini standing as the best approach. Furthermore, a subsequent study statistically analysing the results in datasets with $6$ or more classes, where the ordinality of the output variable is more pronounced, proves significant differences in favour of the ordinal approaches. Remarkably, OGini significantly outperforms both Gini and IG in MAE and QWK.
All these results, source code, implementation of the considered splitting criteria, and instructions on reproducing the experiments are available in the associated website [^3] https://www.uco.es/ayrna/grupos/ordinal-trees.
We recognise the potential for further development of effective ordinal splitting criteria. In this sense, we also provide a framework that eases the testing and validation of new DT methods and splitting criteria in a plug-and-play way. Hence, the unified mathematical notation, code, and experimental study presented in this work offer additional opportunities to explore this field, serving as a baseline for future developments.
A potential future line of research is to study the application of these splitting criteria in ensembles of trees, where the ordinality can be further exploited employing ordinal ensemble techniques such as the ordinal forest proposed in [54]. We also would like to extend the ordinal problems archive used for comparison, as this will facilitate the validation of new proposals. We greatly appreciate any contributions to this archive.
Acknowledgments
The present study has been supported by the “Agencia Estatal de Investigación (España)” (grant ref.: PID2023-150663NB-C22 / AEI / 10.13039 / 501100011033), by the European Commission, AgriFoodTEF (grant ref.: DIGITAL-2022-CLOUD-AI-02, 101100622), by the Secretary of State for Digitalization and Artificial Intelligence ENIA International Chair (grant ref.: TSI-100921-2023-3), and by the University of Córdoba and Junta de Andalucía (grant ref.: PP2F_L1_15). R. Ayllón-Gavilán has been supported by the “Instituto de Salud Carlos III” (ISCIII) and EU (grant ref.: FI23/00163).
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Author contributions: CRediT
R. Ayllón-Gavilán: Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Writing - Original draft. F. J. Martínez-Estudillo: Conceptualization, Formal analysis, Investigation, Methodology, Writing - Original draft, Writing - review and editing. D. Guijo-Rubio: Resources, Validation, Visualization, Writing - review and editing. P. A. Gutiérrez: Data curation, Funding acquisition, Project administration, Supervision, Writing - review and editing. C. Hervás-Martínez: Conceptualization, Formal analysis, Resources, Supervision, Writing - review and editing.